OpenStudy (anonymous):

Can someone please explain to me how to solve this problem? 6-4(3-5)^2+30÷5 Like I know the answer but I don't know the steps to solving it :/

OpenStudy (zehanz):

You have to respect the order of operations: first calculate powers, then multiply and divide (no priority for either of them) then add and subtract (same here). Brackets can change this: you have to calculate what is between them first. So: 3-5=-2. we then have 6-4*(-2)^2+30/5. Powers first: (-2)^2=(-2)*(-2)=4, so: 6-4*(-2)^2+30/5 = 6-4*4+30/5. Now multiply and divide: 6-4*4+30/5=6-16+6. Now add and subtract (from left to right, no priority for + or -) Answer: -4

OpenStudy (zehanz):

YW!

OpenStudy (anonymous):

thanks!! :)